### v1

# One Diffusion to Generate Them All

Abstract
We introduce OneDiffusion, a versatile, large-scale diffusion model that seamlessly supports bidirectional image synthesis and understanding across diverse tasks. It enables conditional generation from inputs such as text, depth, pose, layout, and semantic maps, while also handling tasks like image deblurring, upscaling, and reverse processes such as depth estimation and segmentation. Additionally, OneDiffusion allows for multi-view generation, camera pose estimation, and instant personalization using sequential image inputs. Our model takes a straightforward yet effective approach by treating all tasks as frame sequences with varying noise scales during training, allowing any frame to act as a conditioning image at inference time. Our unified training framework removes the need for specialized architectures, supports scalable multi-task training, and adapts smoothly to any resolution, enhancing both generalization and scalability. Experimental results demonstrate competitive performance across tasks in both generation and prediction such as text-to-image, multiview generation, ID preservation, depth estimation and camera pose estimation despite relatively small training dataset. Our code and checkpoint are freely available at https://github.com/lehduong/OneDiffusion.
Figure 1: OneDiffusion is a unified diffusion model designed for both image synthesis and understanding across diverse tasks. It supports text-to-image generation (red box), conditional image generation from input images (orange box) and it’s reverse task Image understanding (orange box). It can also perform ID customization (blue box), and multi-view generation (purple box) with arbitrary number of input and output images.
1 Introduction
Diffusion models, particularly in text-to-image (T2I) generation, have recently shown remarkable results. Models such as DALL-E [43], Imagen [43], and Stable Diffusion [47, 41, 15] have established new benchmarks for generating high-quality, photorealistic images from text prompts. Additionally, recent studies have demonstrated the effectiveness of diffusion models in various other computer vision tasks, such as depth estimation [22] or optical flow estimation [49, 37], etc. However, despite these advancements, diffusion models are typically trained individually for either T2I generation or specific tasks.
In contrast, large language models (LLMs) (e.g. GPT-4 [1]) have demonstrated their ability to function as universal models. They can perform a wide range of tasks across different domains without the need for task-specific modules, and can effectively handle tasks they haven’t been explicitly trained on zero-shot. This universality has been immensely valuable; it has dramatically simplified using, training and scaling these models, and ultimately lead to better performance. This incentivizes us to ask whether diffusion models can become universal in a similar way.
Designing a unified architecture for diverse image synthesis tasks presents significant challenges. Current methods often depend on external add-ons to handle new tasks. For example, ControlNet [68] or T2I-Adapter [39] require a specialized module to encode the conditional inputs, and personalization models typically require encoding the identity through a pretrained facial recognition network and adding auxiliary losses to preserve identity [63, 58, 21]. Additionally, tasks vary widely in input requirements. For instance, multi-view generation alone requires handling arbitrary input-output view combinations, posed or unposed images, and camera pose conditioning [18, 50, 57, 34, 24], while image understanding tasks require diverse outputs such as depth, pose, or segmentation. Finally, existing training recipes are often tightly tuned to particular tasks and therefore cannot be relied on to generalize between tasks.
In this work, we present OneDiffusion – a unified diffusion model that seamlessly supports bidirectional image synthesis and understanding across diverse tasks. Our approach enables a single model to perform multiple tasks without the need for external losses and add-ons. Inspired by recent advances in diffusion models for sequential data [69, 48, 7], we model all conditions and target images as a sequence of “views” with varying noise levels during training. At inference, any of the views can be used as a conditional input, or set to noise and then used to generate an output image. Conditioning text can also be changed to define the task, and specify additional conditioning details (e.g. camera pose). The simple, but flexible, framework allows our model to support many kinds image generation and image understanding tasks with a unified architecture and training objective.
To train our model we create the One-Gen dataset, which integrates high-quality data across a variety of sources. This includes standard T2I data along with synthetic outputs from state-of-the-art models to support a range of tasks such as depth estimation, segmentation, pose estimation, etc. Our dataset also incorporates data for ID customization and multiview generation, providing diverse conditioning setups. The One-Gen dataset enables scalable joint training across multiple tasks, flexible conditioning options, improved generalization, and serves as a solid foundation for our model.
To demonstrate how general-purpose our training algorithm is, we train OneDiffusion completely from scratch. First, we train on text-to-image to equip the model with general image synthesis abilities, then on One-Gen to learn the full set of tasks. Our final model has 2.8 billion parameters and is equipped with a diverse set of skill, shown in Figure 1. The model also adapts naturally to various resolutions, enabling zero-shot high-resolution generation for sequence tasks even when such resolutions were not encountered during training.
We evaluate OneDiffusion on a diverse set of tasks for both generative and predictive tasks. On T2I, OneDiffusion efficiently generates high-quality images while utilizing fewer number of parameters. In the multiview generation task, OneDiffusion demonstrates performance comparable to state-of-the-art methods that are specifically designed and exclusively trained for this purpose. On depth estimation, OneDiffusion exhibits performance that is competitive with SoTA diffusion-based depth estimation baselines. We also show OneDiffusion supports novel conditioning setups, such as text-to-multi-view and image-to-multi-view. For high-variability tasks like face identification from a single image, the model is capable of generating multiple consistent images featuring flexible expressions and poses, demonstrating strong generalization to unseen domains.
2 Related work
Diffusion models for generative task.
Recent advancements in diffusion models have greatly improved image generation capabilities, with models like Stable Diffusion [47, 41, 15, 4, 3, 72] setting new standards in text-to-image synthesis. Building on this foundation, controllable diffusion models such as ControlNet [68] and T2I-Adapter [39] extend functionality by enabling fine-grained control through auxiliary inputs like edge maps, depth maps, or human poses. Meanwhile, instruct-Pix2Pix [5] introduces natural language-guided image editing, making these tools more user-friendly. For personalized applications, identity-focused models, including IP-Adapter [63], InstantID [58], PhotoMaker [27], and PuLiD [21], personalize generation by conditioning on reference images. Meanwhile, in multi-view generation, recent methods [50, 57, 34, 18], employ camera ray embeddings or 3D geometry to achieve consistent viewpoints. Together, these innovations showcase the versatility of diffusion models in delivering controllable, personalized, and multi-perspective image synthesis.
Diffusion models for predictive tasks.
Beyond image generation and manipulation, diffusion models have also proven effective for predictive tasks within computer vision. Marigold [22] fine-tunes the Stable Diffusion model [47] to perform monocular depth estimation, demonstrating the adaptability of diffusion models for prediction-based applications. Furthermore, diffusion models have been utilized for optical flow estimation, as shown in the works of Saxena et al. [49] and Luo et al. [37], where the models predict pixel-level motion between consecutive frames. Additionally, Li et al. [26] trained a diffusion model for open-vocabulary semantic segmentation, showcasing the potential of these models for more complex vision tasks. Prior works have attempt to unify diffusion model for predictive tasks [19, 17]. These studies show that diffusion models are not only useful for generating images but also highly effective for various predictive tasks in computer vision.
Unified diffusion models.
Several attempts have been made to unify diffusion model for different type of controls [70, 42, 60]. However, they are limited to utilization of multiple image conditions. These models usually requires to design complicated adapters for different conditions. [35, 36, 55, 71] propose unified models for language and images. Concurrently, [59] propose finetuning multimodal large language model with diffusion objective on diverse tasks like text-to-image, editing, and subject-driven generation etc. In contrast, our model distinguishes itself by leveraging bidirectional capabilities of diffusion models and addressing a wide range of diverse tasks.
Figure 2: Illustration of training and inference pipeline for OneDiffusion. We encode the desired task for each sample via a special task token. During training we independently sample different diffusion timesteps for each view and add noise to them accordingly. In inference, we replace input image(s) with Gaussian noises while setting timesteps of conditions to 00.
3 Methodology
3.1 Flow matching for generative modeling
Flow matching [30, 33, 2] is a framework for training continuous-time generative models by learning a time-dependent vector field that transports between two probability distributions. More specifically, a time-dependent vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ governs the transformation from a base distribution $p_{0}$ to the target distribution $p_{1}\approx q$ through an ODE $dx=u_{t}(x)dt$.
The solution of this ODE is a flow $\phi_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with initial condition $\phi_{0}(x)=x$, and this flow characterizes a push-forward operation $p_{t}=[\phi_{t}]_{\#}p_{0}$, in which $p_{t}$ is the density of samples $x\sim p_{0}$ transported by $u$ from time 00 to time $t$. The goal is approximate this ODE using a learned time-dependent vector field parameterized as a neural network $v_{\theta}(t,x)$. Due to the intractable nature of $u_{t}$, [30] proposed to learn $v_{\theta}(t,x)$ using the conditional flow matching (CFM) objective:
$\mathcal{L}_{\mathrm{CFM}}(\theta):=\mathbb{E}_{t,q(z),p_{t}(x\mid z)}\left\|v_{\theta}(t,x)-u_{t}(x|z)\right\|^{2}$ (1)
This objective is equivalent to the original flow matching objective, and only requires the samples from the target distribution and a suitable conditional probability path.
3.2 Proposed Approach
Objective.
We cast the problem of image generation with multimodal conditions as sequential modeling. Inspired by previous work on diffusion model for sequential data [69, 48, 7], we jointly model all conditions and target images as a sequence of “views”. Note that the number of views $N$ is determined by tasks. Particularly, $N=1$ for text-to-image tasks, $N=2$ for image-to-image translation such depth/pose/image editing etc, $N>2$ for multiview generation or ID customization.
Mathematically, let $N$ views $\{\mathbf{x}_{i}\}_{i=1}^{N}\in\mathbb{R}^{H\times W\times D}$ be sampled from a training dataset $q(\mathbf{x}_{1},...,\mathbf{x}_{N})$. Given time variables $t_{i}$, our goal is to learn a function $v_{\theta}(t_{1},...,t_{N},\mathbf{x}_{1},...,\mathbf{x}_{N}):[0,1]^{N}\times\mathbb{R}^{N\times H\times W\times D}\rightarrow\mathbb{R}^{H\times W\times N\times D}$. Intuitively, $v_{\theta}$ serves as a generalized time-dependent vector field where each input $\mathbf{x}_{i}$ paired with its respective time variable $t_{i}$.
Learning $v_{\theta}$ enables arbitrary conditional generation, where any subset of views can be selected as conditions to generate the remaining views, as explained below. This setup allows us to dynamically configure the generation process, supporting flexible applications across a range of generative tasks.
Training.
Our training pipeline is visualized on the left side of Figure 2. At each training step, we independently sample $t_{i}\sim\text{LogNorm}(0,1)$ [15] and Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,I)$. This results in different noise levels for each views. We apply an interpolation-based forward process:
$\mathbf{x}_{i}^{t_{i}}=\alpha_{t_{i}}\mathbf{x}_{i}+\beta_{t_{i}}\epsilon_{i}$ (2)
where $\alpha_{t}$ and $\beta_{t}$ satisfy the boundary conditions $\alpha_{0}=0,\alpha_{1}=1$ and $\beta_{0}=1,\beta_{1}=0$. Similar to [72], we adopt the linear interpolation schedule:
$\mathbf{x}_{i}^{t_{i}}=t_{i}\mathbf{x}_{i}+(1-t_{i})\epsilon_{i}$ (3)
the corresponding velocity field $u_{i}$ for each view $\mathbf{x}_{i}$ is:
$u_{i}(t_{i},\mathbf{x}_{i})=\mathbf{x}_{i}-\epsilon_{i}$ (4)
with the aggregated target as $u=(\mathbf{x}_{1}-\epsilon_{1},...,\mathbf{x}_{N}-\epsilon_{N})\in\mathbb{R}^{N\times H\times W\times D}$, our training loss is the joint flow-matching objective:
$\mathcal{L}(\theta)=\mathbb{E}\left[\|v_{\theta}(t_{1},...,t_{N},\mathbf{x}_{1},...,\mathbf{x}_{N})-u\|^{2}\right]$ (5)
This flow matching objective [30, 33, 2] guides the model to learn the optimal velocity field $v_{\theta}$ by minimizing the difference from the target velocity field $u$.
Inference.
Our framework allows for both joint sampling and conditional sampling across any chosen set of views. In details, we define the target views we want to sample as $\mathbf{x}_{K}=\{\mathbf{x}_{i}\}_{i\in K}$, and the set of conditional views as $\mathbf{x}_{\setminus K}=\{\mathbf{x}_{i}\}_{i\notin K}$. To perform conditional sampling, we start by initializing the target views $\mathbf{x}_{K}$ as Gaussian noise. At each timestep $t$, we compute the corresponding time-dependent vector field $v_{\theta}^{K}(t,\mathbf{x}|\bar{\mathbf{x}}_{\setminus K})$ by fixing the conditional views to their known values $\mathbf{x}_{\setminus K}=\bar{\mathbf{x}}$ and setting their time variables to zero $t_{\setminus K}=0$; while keeping the time variables of the target views as $t_{K}=t$:
$v_{\theta}^{K}(t,\mathbf{x}|\mathbf{x}_{\setminus K}=\bar{\mathbf{x}})=v_{\theta}(t_{K}=t,t_{\setminus K}=0,\mathbf{x}_{K}=\mathbf{x},\mathbf{x}_{\setminus K}=\bar{\mathbf{x}})$ (6)
Note that unlike $v_{\theta}$, $v_{\theta}^{K}$ is a valid time dependent vector field, as all the views in $K$ now has the same $t$. Thus, by integrating this vector field using an ordinary differential equation (ODE) solver, we can generate the conditional samples we are interested in. We illustration the inference on the right side of Figure 2.
Figure 3: High-resolution samples from text of our OneDiffusion model, showcasing its capabilities in precise prompt adherence, attention to fine details, and high image quality across a wide variety of styles.
3.3 Implementation Details
Model architecture.
We adopt the Next-DiT architecture [72] in our model. By leveraging a full transformer-based architecture, our model can work with different numbers of views $N$. We independently encode each frame i.e. images and conditions as latent $\mathbf{z}\in\mathbb{R}^{N\times H\times W\times C}$ with a VAE tokenizer [15] and concatenate them in $N$ dimension. With flexible $N$, our approach establishes a universal framework which supports diverse input-modality with variable length. Following [72], we also apply 3D RoPE [53] for positional encoding, enabling generalization to different resolutions and aspect ratios.
Text-to-Image (1 views).
With only a single “view”, training and inference follow the same process as standard text-to-image diffusion models. We prepend the task label [[text2image]] to the caption to specify the task.
Image-to-Image (2 views).
We set the first view as the target image and the second as the conditioning input. During inference, we can use one or both views for generation, and the model is trained to produce the target image. For tasks like bounding box or semantic map generation, we add the hexadecimal color code and class label to the prompt. For instance, to segment a mouse with a yellow mask, the prompt is: [[semantic2image]] <#FFFF00 yellow mask: mouse> photo of a … Further details are provided in the appendix.
ID Customization (2-4 views).
We sample images of the same individual across views, concatenating captions for each input image and using a token [[imgX]] to denote each image. We also prepend the task label [[faceid]] to the captions. At inference, we can condition on an arbitrary number of images and generate multiple outputs, leading to more consistent results.
Multiview Generation (4-12 views).
Inspired by [18], we use Plücker ray embeddings to represent camera poses. For each image patch, we calculate Plücker coordinates as $\bm{r}=(\bm{o}\times\bm{d},\bm{d})$ using its ray origin $\bm{o}$ and direction $\bm{d}$. The result embedding has dimensions $H/8\times W/8\times 6$, matching the spatial size of the latent, and is replicated across channels to form a $16$ channel embedding. Unlike [18], we treat ray embedding as a independent “view” following image latents as a unified sequence rather than concatenating by channels. This design allows flexible denoising, enabling multi-view image generation conditioned on camera poses or sampling ray embeddings to predict poses from image conditions, similar to RayDiffusion [67]. We scale the ray embeddings to have unit variance, as in [47].
Figure 4: Illustration of our model capability to generate HED, depth, human pose, semantic mask, and bounding box from input image. For semantic segmentation, we segment the sword (highlighted in yellow) and the moon (highlighted in cyan) the first example, while segmenting road (yellow), sky (cyan) in the second. For object detection, We localize the head and moon (both highlighted in cyan). Leveraging these conditions, we can reverse the process to recreate a variant of the input image based on the same caption. Additionally, we can edit the image by modifying specific elements, such as replacing the moon with Saturn (last example).
As with other tasks, we prepend the task label [[multiview]] to the caption. During inference, we can substitute images or Plücker ray embeddings with Gaussian noise for multi-view generation and camera pose estimation, respectively.
Training details
Our model is trained from scratch using a flow-matching objective. Similar to prior works [15, 8], we use a three stage training recipe. In the first stage, we pretrained the text-to-image model with resolution of $256^{2}$ (500K steps) and $512^{2}$ (500K steps). In the second stage, we continue training on a mixed of tasks, using $512^{2}$ for T2I and $256^{2}$ for other tasks, for a total of 1M steps. Finally, in the last stage, we finetune the model at a high resolution of ($1024$) for T2I. For ID customization fine-tuning, we use 2-5 views. For fewer views (2-3), we apply a resolution of $512^{2}$, while for more views, we use $256^{2}$ resolution.
During training, we use an in-batch sampling strategy at each stage, sampling tasks (T2I, Image-to-Image, ID customization, and multiview generation) with equal probability. The noise scheduler’s shift value is set to 3, as suggested in [15]. We use AdamW optimizer with learning rate $\eta=0.0005$. Training is performed on a TPU v3-256 pod with a global batch size of 256 in the first two phases, and the final fine-tuning stage is completed on 64 H100 GPUs using the same configuration.
4 One-Gen Datasets
Text-to-Image
We leverage both public and internal (synthetic) datasets. The public datasets including: PixelProse [52], Unsplash, Coyo [6], JourneyDB [40]. Additionally, we use a 10M internal synthetic dataset consisting of images re-captioned with LLaVA-NeXT [31] and Molmo [11]. The length of the text description for each image varies from $100$ to $150$ words. When an original prompt is available, we use both the LLM-generated caption and the original caption.
Image-to-Image
For simpler tasks like deblurring, inpainting, image generation from canny edge, or upscaling, we use a $1M$-sample subset of our synthetic data and apply the relevant pre-processor for each image to create an input condition for it. For more complex tasks, we create a synthetic dataset from outputs generated by Midjourney, Stable Diffusion [15], and Flux-dev following the process outlined below:
  • Semantic Map and Detection For each image, we use the LLaVA-NeXT [31] model to identify entities or subjects (e.g., person, shirt, dog, building), with a maximum of 10 entities per image. Based on these subject names from LLaVA-Next, we perform semantic segmentation using SAM [23] and extract bounding boxes. Each class is assigned a random color from a predefined list. This dataset contains 350K triplets consisting of a semantic map, bounding box, and the original image.
  • Depth Map We generate the depth dataset by applying DepthAnything-v2 [62] to 500K images sampled from various datasets, including both real and synthetic images. Additionally, we caption 40K images from Hypersim dataset [46] with LLaVA-NeXT and incorporate these into the training set.
  • Human Poses We collect a different subset with 50K images, primarily featuring human for pose conditioning. We use YOLOv5 to detect the bounding boxes for region of interests and apply ViTPose [61] for pose estimation.
ID Customization
We collect a dataset of celebrities and characters from games and movies by from publicly available images. After filtering to ensure each subject has at least four images and removing NSFW content, the dataset includes approximately 60K subjects and a total of 1.3M images. We caption these images using the LLaVA-NeXT.
Multiview Generation
We use the DL3DV-10K dataset [29], Objaverse [10], CO3D [45]. For Objaverse dataset, we utilize the 80K filtered split from LGM [54] and caption provided by Cap3D [38]. In the DL3DV dataset, we sample an image from each scene and caption it using LLaVA-Next. For CO3D, we exclude captions and include only the task token in the text input.
Figure 5: Illustration of the multiview generation with single input image. We equally slice the azimuth in range of $[-45,60]$ and elevation in range of $[-15,45]$ for the left scenes. For the right scene, the azimuth range is set to $[0;360]$ and elevation range is set to $[-15;15]$.
5 Experiments
In this section, we evaluate our OneDiffusion model on broad range of image generation and understanding tasks. We do not perform task-specific finetuning in any results. Details about additional qualitive exampels are in Appendix.
5.1 Text-to-Image
Qualitative results of OneDiffusion for text-to-image task is illustrated in Figure 3. Thanks to the diversity of our One-Gen dataset, the model can handle various art styles, spanning both artistic and photorealistic designs.
Following previous works [15], we evaluated the text-to-image capabilities of our model on GenEval benchmark [20]. For each prompt, we generate 4 images using Euler solver with $100$ steps and guidance scale of $5$. The results for OneDiffusion, along with those of baseline models, are presented in Table 1. Our model demonstrates strong performance compared to similarly sized baselines, excelling in multitasking capabilities despite being trained on a relatively smaller dataset.This performance is largely attributed to the diversity of the dataset and the comprehensive captions provided for each sample.
Methods Params (B) # Data (M) GenEval $\uparrow$
LUMINA-Next [72] 2.0 14 0.46
PixArt-$\Sigma$ [9] 0.6 33 0.54
SDXL [41] 2.6 0.55
PlayGroundv2.5 [25] 2.6 0.56
IF-XL 5.5 1200 0.61
SD3-medium [15] 2.0 1000 0.62
Hunyuan-DiT [28] 1.5 0.63
DALLE3 0.67
FLUX-dev 12.0 0.67
FLUX-schnell 12.0 0.71
OneDiffusion 2.8 75 0.65
Table 1: Comparison of text-to-image performance on the GenEval benchmark at a resolution of 1024 $\times$ 1024.
5.2 Controllable Image generation
We show the experiment with image-to-image translation using various source domains, including HED, depth map, human pose, semantic map, bounding boxes. We report the qualitative results in Figure 4 and Figure 19 in appendix. The results demonstrate that OneDiffusion effectively aligns with the conditioning image across a wide range of input conditions during the generation process, leveraging pure attention mechanisms and supplementary information from captions.
5.3 Multiview Generation
Figure 6: Illustration of ID customization using reference images. Unlike prior methods that rely on face embeddings and often fail to generalize, our model demonstrates superior generalization. It effectively adjusts facial expressions and gaze directions (first row), changes viewpoints (second row), and even customizes non-human IDs (third row). All results in the third row are generated from a single reference image, while InstantID fails as its face detector cannot detect faces in the input.
Model Condition PSNR $\uparrow$
Zero123 [32] 1-view 18.51
Zero123-XL [12] 1-view 18.93
1-view 20.24
EscherNet [24] 2-view 22.91
3-view 24.09
1-view 19.01
2-view (unknown poses) 19.83
OneDiffusion 2-view (known poses) 20.22
3-view (unknown poses) 20.64
3-view (known poses) 21.79
Table 2: Comparison of NVS metrics across different number of condition view settings. Increasing the number of condition views improves the reconstruction quality.
We evaluate the multiview generation capabilities of our method using the Google Scanned Object dataset [13]. Table 2 presents a comparison of our approach (OneDiffusion) with state-of-the-art methods, including Zero123 [32], Zero123-XL [12], and EscherNet [15]. It is important to note that these models are specifically designed and exclusively trained for the multiview generation task. In contrast, our model supports a variable number of conditional inputs and, due to its flexible denoising framework, can incorporate additional conditional views even when their camera poses are unknown.
As shown in Table 2, OneDiffusion consistently outperforms Zero123 and Zero123-XL in the 1-view condition. Additionally, OneDiffusion demonstrates strong performance even when camera poses are unknown. For instance, in the 2-view condition with unknown poses, OneDiffusion achieves a PSNR of 19.83, which is only marginally lower than the PSNR of 20.22 obtained when camera poses are known. Similarly, in the 3-view condition, OneDiffusion achieves PSNR values of 20.64 and 21.79 for unknown and known poses, respectively. These results highlight the adaptability and effectiveness of our method across varying input conditions, demonstrating its potential for practical applications in multiview generation.
Figure 5 presents two qualitative examples of our multiview generation from a single front-view image. Our model successfully generates consistent views across varying azimuths and elevations. For additional visualizations, please refer to Figures 10 and 11 in Appendix. As a result of the flexibility of our framework, we can also perform text-to-multiview directly by “masking” all images and only inputing the camera poses as in Figure 12 in Appendix.
5.4 ID Customization
We further evaluate OneDiffusion on ID customization tasks, which involve using one or multiple ID images as inputs for personalized generation. To assess performance, we compare with STOA methods, including InstantID [58], PuLID [21], and PhotoMaker [27], using both qualitative and quantitative analyses. Our evaluation extends beyond the standard benchmark (unsplash-50 [16]) to test generalization on ID customization tasks, such as varying expressions, viewpoints, and even non-human images.
Figure 6 illustrates examples of altering facial expressions and gaze directions (first row), changing viewpoints (second row), and customizing non-human IDs (third row). Our method achieves success in these tasks, where all other methods fail. Unlike previous approaches that rely on face embeddings and primarily “replicate” the original face, OneDiffusion employs attention mechanisms between images and text conditions. This enables flexible end-to-end training and generates more expressive outputs, making our method suitable for a wider range of applications. Intuitively, the mechanism that ensures consistent multiview generation also proves effective for manipulating camera angles in ID customization, highlighting its adaptability across related applications. Additional visualizations are provided in Figure 13 and 14.
Method ID $\uparrow$ CLIP-T $\uparrow$
PhotoMaker [27] 0.193 27.38
InstantID [58] 0.648 26.41
PuLID [21] 0.654 31.23
Ours 0.283 26.80
Table 3: Quantitative results on Unsplash-50.
We also present the quantitative results on the Unsplash-50 [16] benchmark in Table 3. This benchmark focuses solely on style changes and re-contextualization, where PuLID [21] demonstrates strong performance by leveraging embeddings from ID encoder networks trained on human faces for discrimination tasks. While this approach effectively preserves the identity traits of input images, it faces significant limitations when handling more complex face manipulations.
5.5 Depth Estimation
Figure 7: Qualitative comparison between a diffusion-based depth estimation - Marigold [22] and our methods.
Method NYUv2 DIODE
[…]
DiverseDepth [64] 11.7 87.5 37.6 63.1
MiDaS [44] 11.1 88.5 33.2 71.5
DPT [44] 9.8 90.3 18.2 75.8
LeReS [65] 9.0 91.6 27.1 76.6
Omnidata [14] 7.4 94.5 33.9 74.2
HDN [66] 6.9 94.8 24.6 78.0
Marigold [22] 6.0 95.9 31.0 77.2
DepthAnything-2 [62] 4.6 97.7 27.1 74.8
Ours 6.8 95.2 29.4 75.2
Table 4: Comparison of depth estimation methods on NYUv2 and DIODE datasets.
For image understanding tasks, we evaluate our model’s performance on monocular depth estimation using standard benchmarks: NYUv2 [51] and DIODE [56]. We report the quantitative results in Table 4.Quantitative results are presented in Table 4, where our model achieves competitive performance compared to baselines that leverage pretrained text-to-image diffusion models, such as Marigold [22]. However, as shown in Figure 7, our model demonstrates greater robustness than diffusion-based depth estimators like Marigold. Specifically, it excels in handling open-world images, including paintings, hazy weather, and unconventional textures. For further qualitative comparisons, refer to Figures 15 and 16.
5.6 Camera Pose Estimation
We evaluate our model on camera pose estimation using the Google Scanned Object dataset [13]. For this task, we use six rendered images of each synthetic object and estimate the camera poses by denoising the corresponding ray embeddings. Following RayDiffusion [67], we apply least squares optimization to estimate the camera centers and rotations. The camera center accuracy, measured with a threshold of 0.3, is reported in Table 5.
Figure 8 provides a qualitative comparison between our model and RayDiffusion. RayDiffusion consistently predicts camera poses in the upper hemisphere due to the bias in its training data, such as CO3D, which predominantly features upper-hemisphere views. In contrast, thanks to the diversity of our large-scale training dataset, OneDiffusion achieves higher accuracy and avoids this limitation.
Figure 8: Qualitative comparison between RayDiffusion and OneDiffusion on GSO dataset. OneDiffusion yields better prediction.
Method Accuracy
RayDiffusion [67] 0.20
OneDiffusion 0.32
Table 5: Comparison of zero-shot camera pose estimation methods on the GSO dataset, evaluated by Camera Center Accuracy at a threshold of 0.3.
5.7 Other Tasks
Since directly extracting masks, bounding boxes, and keypoints from raw output images is not straightforward, we provide several qualitative results for human pose estimation and semantic segmentation on COCO dataset in Figures 17 and 18 in the appendix, respectively. Since our model does not distinguish tasks of conditions and images during training, its performance on understanding tasks serves as an additional indicator of the model. We leave further exploration of this aspect for future work.
6 Conclusion
Our experiments demonstrate that OneDiffusion achieves impressive results across a variety of tasks, including conditional T2I generation, depth estimation, open vocabulary semantic segmentation, pose estimation, multi-view generation, ID customization and camera pose estimation. We believe this work advances the capabilities of diffusion models, providing a versatile and scalable solution comparable to the flexibility offered by large language models. This represents a significant step toward developing a general-purpose vision model that can serve as the backbone for a wide variety of applications.
Supplementary Material
7 Additional Qualitative results
ID Customization.
We report ours results for ID Customization tasks in Figure 13 and Figure 14. It can be observed from Figure 13 that OneDiffusion well preserve the identity of a person with a single input and highly manipulatable. It can re-contextualize the image (as in first prompt), change the style from realistic photo to Pixar style (second prompt) and modify the medium to watercolor painting (third prompt).
Moreover, our approach does not relying on face embedding as in previous works [21, 27, 58] making it highly versatile. As illustrated in Figure 14, our model can preserver highly details, intricate structure as armor of the person in $4^{th}$ row. OneDiffusion can also work with non-human subject as the Gundam robot in $3^{rd}$ row. The model performs well with other style than photorealistic input such as anime ($1^{st}$ row), 3D figure ($2^{nd}$ row), cartoon ($5^{th}$ row). Our model is highly editable where we can control the style, human pose, camera angle, expression.
Multiview generation
We report additional results for multiview generation in Figure 10 and Figure 11. The generation process is as follow: we set the azimuth ranges to $[-0.45,0.6]$ and elevation ranges to $[-15,45]$, except for the last row of Figure 11. Then we equally slice these ranges to $80$ views. We first generate 3 anchor views from the input image and independently synthesize subsequent images based on input image and the nearest anchor. For each generation batch, we generate $3$ novels view and condition on $2$ images. We report views with index in $[0,10,20,\cdots,70]$ in below figures.
OneDiffusion is capable of generating photorealistic results of arbitrary objects or scenes from any number of input views either realistic captured ($2^{nd},3^{rd}$ row of Figure 11) or synthesized images (Figure 10). Our model works best for camera trajectory covering front views of a scene.
As mentioned ealier, OneDiffusion can also generate consistent multiview images from pure text and without any input images. Specifically, we simply input the azimuth and elevation as input for camera poses and generate all images from Gaussian noises as in Figure 12.
Depth estimation
We provide additional qualitative results of OneDiffusion and compare it with Marigold-LCM [22] and DepthAnything-v2 [62] in Figure 15 and 16. We can see that our model estimator is more robust than Marigold on open-world test suits and is highly correlated with output of DepthAnything model.
Human Pose estimation
We report additional results for pose estimation on COCO dataset in Figure 17. It can be observed that our model can predict multiple people in an image without relying on object detector models.
Semantic Segmentation
We report qualitative results of semantic segmentation on COCO dataset in Figure 18. Unlike previous models [42, 70], our semantic-to-image and vice verse does not enforce hard association between colors and the target classes. We provide the color masks and class name as additional input in caption.
Figure 9: Distribution of training datasets for all tasks. Segments proportional to sampling rates. The inner section shows the super-category of target tasks, it can be observed that we train the model with equal budget for text-to-image, image-to-image and multiview generation. The outer section shows datasets used for each super-category.
8 Summary Datasets
We train the model on multiple datasets reported in Section 4 and illustrated in Figure 9. The pie-chart segment each dataset proportional to the sampling rate of it in third stage of the training process. We train the model with equal budget for text-to-image, image-to-image translation ($2$ frames), and multiview generations ($2-6$ frames). Note that we filter and only use a subset of COYO with $11M$ images in our training. Due to the missing samples during download process, the LAION-aesthetic dataset only has $6M$ images. We recaption the LAION-aesthetic dataset with Molmo [11].
Figure 10: Qualitative results of image-to-multiview generation. The left most images are input. We equally slice the azimuth in range of $[-45,60]$ and elevation in range of $[-15,45]$ for all scenes.
Figure 11: Qualitative results of image-to-multiview generation. We equally slice the azimuth in range of $[-45,60]$ and elevation in range of $[-15,45]$ for the first 3 scenes. For the last scene, the azimuth range is set to $[0;360]$ and elevation range is set to $[-15;15]$.
Figure 12: Qualitative results of text-to-multiview generation. The azimuth and elevation of left to right columns are $[0,30,60,90]$ and $[0,10,20,30]$, respectively. We use following prefix for all prompts to improve the quality and realism of generated images: “photorealistic, masterpiece, highly detail, score_9, score_8_up”.
Figure 13: Qualitative results of OneDiffusion for (single reference) ID Customization task with photo of human faces. The left most images are input, target prompts for left to right columns are: 1) “Photo of a man/woman wearing suit at Shibuya at night. He/She is looking at the camera”, 2) “pixarstyle, cartoon, a person in pixar style sitting on a crowded street”, 3) “watercolor drawing of a man/woman with Space Needle in background
Figure 14: Qualitative results of OneDiffusion for (single reference) ID Customization task with photo of of non-human subjects or cartoon style input. OneDiffusion is highly versatile and can produce good results for all kind of input and not limited to photorealistic human images. Since we rely on attention, the model can attend to the condition view and preserve intricate details and is not limited by any bottleneck e.g. latent representation.
Figure 15: Qualitative comparison for depth estimation between OneDiffusion, Marigold [22] and DepthAnything-v2 [62]
Figure 16: Qualitative comparison for depth estimation between OneDiffusion, Marigold [22] and DepthAnything-v2 [62]
Figure 17: Qualitative examples of human pose estimation on COCO datasets.
Figure 18: Qualitative examples of semantic segmentation on COCO datasets. The target class for each image (from left to right, from top to bottom) are (sheep, grass, mountain, sky), (apple, person, building), (vase, flower, ), (dog, book, sheet), (umbrella, person, building, gate), (boat, dock, drum).
Figure 19: Illustration of our model capability to generate semantic mask, detection, human pose, depth, and canny edge from input image. For semantic segmentation, we segment the flower (highlighted in yellow) and the rock (highlighted in green). For object detection, We localize the backpack (highlighted in yellow) and butterfly (highlighted in cyan). Leveraging these conditions, we can reverse the process to recreate a variant of the input image based on the same caption.